Graph the linear function by hand. Identify the slope and​ y-intercept.

g(x)=7-4x

What is the slope or is it undefined?
Answer #1

g (c) = 7-400 Salution & g(x) = 7-4x g(x) =-4x+7 compare this with glxl = mx + c slope - 4 AOS for y intercept put x=0 in giv